OpenStudy (anonymous):

Write (8a^-6)^-2/3 in simplest form

OpenStudy (anonymous):

you have two jobs one is to find \[8^{-\frac{2}{3}}\] and the other is to multiply \(6\times -\frac{2}{3}\) second job is easy enough do you know how to do the first one?

OpenStudy (anonymous):

\[\huge x^{-n} = \frac{ 1 }{ x^n } ~~\text{and}~~ x^{m/n} = \sqrt[n]{x^m}\]

OpenStudy (anonymous):

oops typo there second is to find \[-6\times -\frac{2}{3}\]

OpenStudy (anonymous):

\[\huge (x^m)^n = x^{m*n}\] this rule as well will help simplify it as well.

OpenStudy (anonymous):

i don't know the first one

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\huge (8a^{-6})^{-2/3} = 8^{-2/3} * a^{-6*\frac{ -2 }{ 3 }}\]

OpenStudy (anonymous):

Do that first.

OpenStudy (anonymous):

k

OpenStudy (anonymous):